THE CONSTRUCTION OF ORTHONORMAL WAVELETS USING SYMBOLIC METHODS AND A

MATRIX ANALYTICAL APPROACH FOR WAVELETS ON THE INTERVAL

FRÉDÉRIC CHYZAK, PETER PAULE, OTMAR SCHERZER, ARMIN SCHOISSWOHL, AND BURKHARD ZIMMERMANN

Abstract. We discuss closed form representations of filter coefficients of wavelets on the real line, half real line and on compact intervals. We show that computer algebra can be applied to perform this task. Moreover, we present a matrix analytical approach that unifies constructions of wavelets on the interval.